The tension in a rope pulling a log is 100N, the mass of the log is 50kg and the frictional force on the log is 20N. What is the acceleration of the log?
1
Expert's answer
2019-04-22T11:07:11-0400

From the second Newton's law:


"ma=F_{net}=T-F_f"

The acceleration of the log:


"a=\frac{T-F_f}{m}=\frac{100-20}{50}=1.6\frac{m}{s^2}"